Give a) the vertex and b) the range of f(x) = -6x2(the 2 is squared) – 12x + 7.
Answer by stanbon(75887)

Give a) the vertex and b) the range of f(x) = -6x2(the 2 is squared) – 12x + 7.
--------------------
f(x) = -6x^2 - 12x + 7
-------
Because of the negative 6 coefficient the parabola opens downward.
So it has a vertex which is at a maximum:
---
The vertex occurs where x = -b/2a = 12/(2(-6)) = -1
----
f(-1) = -6(-1)^2 -12(-1) + 7
f(-1) = -6 +12+7 = 13
---
Vertex (-1,13)
---
So, range is "All Real Numbers less than or equal to 13".
============================
Cheers,
Stan H.
==============